# What are they talking about? Benchmarking Large Language Models for Knowledge-Grounded Discussion Summarization

### v1

@@ Abstract
In this work, we investigate the performance of LLMs on a new task that requires combining discussion with background knowledge for summarization. This aims to address the limitation of outside observer confusion in existing dialogue summarization systems due to their reliance solely on discussion information. To achieve this, we model the task output as background and opinion summaries and define two standardized summarization patterns. To support assessment, we introduce the first benchmark comprising high-quality samples consistently annotated by human experts and propose a novel hierarchical evaluation framework with fine-grained, interpretable metrics. We evaluate LLMs under structured-prompt and self-reflection paradigms. Our findings reveal: (1) LLMs struggle with background summary retrieval, generation, and opinion summary integration. (2) Even top LLMs achieve less than average performance across both patterns. (3) Current LLMs lack adequate self-evaluation and self-correction capabilities for this task.111Our benchmark is available at zhouweixiao/KGDS
@@
Benchmarking Large Language Models for
Knowledge-Grounded Discussion Summarization
Weixiao Zhou Junnan Zhu Gengyao Li Xianfu Cheng Xinnian Liang Feifei Zhai Zhoujun Li††thanks: Corresponding Author State Key Laboratory of Complex & Critical Software Environment, Beihang University State Key Laboratory of Multimodal Artificial Intelligence Systems, Institute of Automation, CAS University of Chinese Academy of Sciences ByteDance Fanyu AI Laboratory wxzhou@buaa.edu.cn junnan.zhu@nlpr.ia.ac.cn
@@ 1 Introduction
Dialogue summarization aims to distill main topics and interactions from a dialogue into a concise and faithful summary Jia et al. (2023); Rennard et al. (2023); Kirstein et al. (2024). Current methods typically rely exclusively on dialogue content across dataset construction Gliwa et al. (2019); Chen et al. (2021); Zhu et al. (2021); Hu et al. (2023), summarization methodologies Lin et al. (2022); Zhou et al. (2023); Tian et al. (2024), and evaluation paradigms Wang et al. (2022); Gao and Wan (2022); Zhu et al. (2023); Gao et al. (2023); Tang et al. (2023, 2024b); Ramprasad et al. (2024). These efforts implicitly assume that "Information in the dialogue is sufficiently complete to support generation of a clearly understandable summary for outside observers."
@@
However, we find this assumption has fundamental limitations and often breaks down, particularly in discussions involving shared background knowledge. Such discussions exhibit two main traits: (1) Information Omission and Implicit Reference: Participants naturally skip mutually known details and frequently use pronouns or phrases to refer to entities or facts within the contextual background. (2) Personal Opinion: Unlike straightforward information interactions in general dialogues, these discussions focus on exchanging viewpoints, with participants expressing personal opinions from various positions or perspectives (see Figure 1). These characteristics make understanding heavily reliant on background knowledge. Consequently, traditional dialogue summarization pattern maintains this dependency, leading to confusion for outsiders unfamiliar with the context, leaving them wondering: "What are they talking about?".
@@
To address this issue, we introduce Knowledge-Grounded Discussion Summarization (KGDS), a novel task specifically designed to combine shared background knowledge with discussion content to produce observer-friendly summaries. We determine that an effective summary must achieve two complementary objectives: (1) bridge the knowledge gap between outsiders and participants by providing necessary background-supporting information; (2) present clear personal opinions of the participants with clarified implicit references. We formalize these requirements by modeling the task output as two components: background summary and opinion summary. The background summary, which retrieves or condenses relevant contextual information, can be either extractive or abstractive, while the opinion summary integrates clarified viewpoints and is inherently abstractive. Thus, we define two standardized summarization patterns: EBS-AOS and ABS-AOS (refer to Figure 1).
To systematically investigate the challenges of KGDS and evaluate LLM capabilities in this task, we make two significant contributions. First, we construct the first KGDS benchmark containing high-quality, multi-domain samples from realistic news discussions through strict consistency-controlled annotation protocols. Each sample comprises structured shared background knowledge, multi-turn human discussion, and expert-annotated evaluation components at multiple granularities. Second, we propose a novel hierarchical evaluation framework with fine-grained and interpretable metrics (as shown in Figure 1). Specifically: (1) For extractive background summary, we utilize retrieval metrics to evaluate recall, precision, and f1-score of supporting paragraphs. (2) For abstractive background summary, inspired by the minimal granularity of atomic facts Liu et al. (2023b) and the excellent human-alignment of LLM-based fact verification Wei et al. (2024), we assess quality via verification of key supporting fact inclusion and nonsupporting fact exclusion, proposing three metrics that multi-dimensionally quantify background coverage, focus, and summary overall performance. (3) For abstractive opinion summary, we evaluate quality by verifying coverage of clear atomic opinions, where each opinion is minimized and implicit references clarified. For uncovered opinions, we define five error types and perform fine-grained error detection. (4) For pattern overall performance, we use geometric mean for evaluation, which aligns with the KGDS workflow and reflects the joint fulfillment rate of both sub-summaries.
We evaluate leading LLMs under single-turn structured-prompt Li et al. (2023) and multi-turn self-reflection Shinn et al. (2023) paradigms. Our analysis reveals several critical insights: (1) LLMs struggle with precision-recall trade-off in background summary retrieval; (2) LLMs frequently miss key facts and retain irrelevant facts in background summary generation; (3) LLMs encounter difficulties with clarifying implicit references in opinion summary integration; (4) Even advanced LLMs achieve less than average performance across both patterns; (5) LLMs lack sufficient self-evaluation and -correction capabilities for KGDS.
Our contributions are as follows: (1) We introduce a novel KGDS task and establish two standardized summarization patterns. (2) We construct the first benchmark for this task and propose a comprehensive hierarchical evaluation framework with interpretable metrics. (3) Our extensive evaluation reveals specific challenges that current LLMs face with KGDS, providing valuable guidance for future improvements in coarse-grained retrieval, fine-grained generation, and knowledge integration.
2 Related Work
Dialogue Summarization.
Dialogue summarization is widely used in practical applications Lin et al. (2021); Hu et al. (2023). Earlier approaches concentrated on feature modeling Chen and Yang (2021); Fang et al. (2022) and pre-training methods Zou et al. (2021); Zhong et al. (2022a). The emergence of LLMs has greatly enhanced dialogue summarization Wang et al. (2023a); Tian et al. (2024); Lu et al. (2025); Zhu et al. (2025), with recent research focusing on evaluating their effectiveness in this field Liu et al. (2024b); Ramprasad et al. (2024); Tang et al. (2024b). However, these studies often overlook the importance of background knowledge.
Knowledge-Grounded Dialogue Generation.
The task involves generating responses that interact with external knowledge sources like documents Wang et al. (2023b, 2024), databases Zhou et al. (2018); Zhang et al. (2020), or retrieval-augmented systems Lewis et al. (2020); Chen et al. (2024). The main distinctions between this task and our KGDS are: (1) Conversation characteristics: The former focuses on knowledge-intensive dialogues, while KGDS deals with background-sparse discussions. (2) Task objectives: The former enhances response quality through knowledge, whereas KGDS utilizes background for complementary summarization.
Summarization Evaluation.
Earlier similarity-based metrics like Rouge Lin (2004), BERTScore Zhang et al. (2019), and MoverScore Zhao et al. (2019) are coarse-grained and often misaligned with human judgment. NLI- and QA-based approaches improve faithfulness evaluation by checking factual consistency Goyal and Durrett (2020); Laban et al. (2022); Zha et al. (2023) or generating relevant questions Wang et al. (2020); Fabbri et al. (2022); Zhong et al. (2022b), but they are limited to faithfulness and require specialized training. Recently, LLM-based reference-free evaluators Liu et al. (2023a); Chen et al. (2023); Wang et al. (2023c); Shen et al. (2023); Fu et al. (2024) have been used, yet they still lack interpretability and fine-grained assessment. To address these, some studies Song et al. (2024a, 2025); Lee et al. (2024); Yang et al. (2024); Scirè et al. (2024); Wan et al. (2024) utilize atomic fact units Liu et al. (2023b); Min et al. (2023) and LLM-based claim verification Tang et al. (2024a); Wei et al. (2024); Song et al. (2024b), achieving more fine-grained evaluation. Among these, FineSurE Song et al. (2024a) is most relevant to our work. However, its metrics lack dimensional consistency (Completeness is measured at the fact-level, while Conciseness is sentence-level), limiting them to isolated sub-aspect evaluation. In contrast, our coverage and focus metrics are both atomic-fact-level, allowing direct F1-score computation for overall quality assessment while maintaining interpretability. Our method can theoretically be adapted for general query-focused summarization, requiring only the maintenance of inner and outer fact sets.
Summarization Benchmark.
Current research often evaluates LLMs across various summarization tasks, including dialogue Tang et al. (2024b), news Goyal et al. (2022); Zhang et al. (2024), multi-document Huang et al. (2024), multilingual Ye et al. (2024), and query-focused Yang et al. (2023); Liu et al. (2024a). Moreover, some studies investigate inherent issues such as hallucinations Belem et al. (2024); Bao et al. (2024) and biases Ravaut et al. (2024); Chhabra et al. (2024). Unlike these, our work uniquely focuses on LLMs within KGDS, evaluating their capabilities in background retrieval, generation, and opinion integration.
@@ 3 Task Formulation
EBS-AOS Pattern.
@@
where is the shared background knowledge of participants, is the discussion grounded in , and is an instruction for the EBS-AOS pattern. is the LLM-driven summarization system. is the EBS, defined as extractive background supporting chunks for from , where chunks represent text sequences with a predefined granularity (e.g., sentences, paragraphs, sections, or window-sliding segments). is the AOS, defined as clear personal opinions of the participants with clarified implicit references.
ABS-AOS Pattern.
@@
where , , , and retain their definitions from Eq. (1). The instruction is specific to the ABS-AOS pattern. is the ABS, defined as abstractive background supporting information for from .
4 Benchmark
@@ 4.1 Preliminary
We establish the benchmark scenario as a two-participant discussion about news content for two reasons: (1) news discussions are highly prevalent in daily communication, and (2) news summarization is a representative subfield of automatic summarization research.
@@
We collect multi-domain (i.e., business, sports, and world) event-rich news articles from Google News222https://news.google.com as shared background knowledge. To ensure data credibility and avoid potential contamination in LLMs Deng et al. (2023); Li and Flanigan (2024), we control the time span of news between Mar and Sep 2024. Additionally, we retain the original paragraph structure and define paragraph-as-chunk as the minimum granularity under the EBS-AOS pattern (see Table 4 in Appendix B for more statistics).
Expert Recruitment.
We recruit four PhD candidates specializing in NLP to engage in our annotation work. Each pair of experts form a collaborative group to conduct the full-process annotation.
4.2 Discussion Construction
@@
The construction follows the sequence of read, understand, then discuss. Initially, for each news article, we require two experts to independently read and thoroughly understand its content. The purpose of this preliminary step is to ensure knowledge consistency. Subsequently, we ask them to engage in a discussion to exchange viewpoints. This process is open-ended, meaning the initiator of the discussion is chosen at random, and the discussion topics can encompass any events, facts, or sub-information contained within the news article (see Table 5 in Appendix B for discussion statistics).
4.3 Annotation and Evaluation for EBS
Back-Supporting Paragraphs Annotation.
For each paragraph in the news article, we ask the two experts involved in the discussion: "Is this paragraph the background-supporting paragraph for discussion?" and require them to answer with either "yes" or "no". We then apply strict consistency control, retaining only paragraphs consistently annotated as "yes" (i.e., back-supporting) or "no" (i.e., back-nonsupporting). For inconsistent annotations (i.e., one "yes" and one "no"), we determine them to be boundary paragraphs and remove from the original news article (see Appendix B for annotation statistics). Our investigation in Section §6.1 shows that LLMs, with an average win rate of , prefer expert-consistently-annotated EBS over their own extractions, highlighting the excellent quality of our annotations.
Evaluation Metrics.
EBS assessment involves matching LLM-extracted paragraph indices. We use information retrieval metrics: Back-Supporting Paragraph (BSP) Recall, Precision, and F1-score (refer to Appendix C for metric details).
4.4 Annotation and Evaluation for ABS
Key Back-Supporting Atomic Facts Annotation.
This process consists of two steps: atomic fact decomposition and expert annotation. We use GPT-4o333Version: gpt-4o-2024-08-06 to decompose back-supporting paragraphs into candidate atomic facts, following the principles of indivisibility, independence, and declarativity (the instruction is provided in Appendix D.2). Next, we ask two experts for each fact: "Is this fact the key background-supporting fact for discussion?" and require them to answer either "yes" or "no". We employ the same consistency control to determine key back-supporting atomic facts (i.e., two "yes") and non-key facts (i.e., one or zero "yes" responses) (refer to Appendix B for annotation statistics).
Back-Nonsupporting Atomic Facts Annotation.
@@ 4.4 Annotation and Evaluation for ABS
This process is automated and includes two steps: atomic fact decomposition and conflicting fact masking. We employ the same approach to decompose back-nonsupporting paragraphs into initial atomic facts. Next, we mask444The masking process is consistent with fact verification, and the conflicting facts are not considered in ABS evaluation. facts that can be inferred from back-supporting paragraphs to obtain back-nonsupporting atomic facts. Conflicts typically occur when identical or similar facts appear in both back-supporting and -nonsupporting paragraphs at the atomic granularity level. For instance, two events have the same timestamp, but one is the background event while the other is not. By masking such facts, we construct non-overlapping back-supporting and -nonsupporting fact sets, avoiding any external verification problems (see Appendix B for annotation statistical data).
Fine-Grained Multi-Dimensional Metrics.
Let denote the LLM-generated ABS, the set of key back-supporting atomic facts, the set of back-nonsupporting atomic facts. The LLM-based fact verification function is defined as:
where and represent whether the fact can be inferred from .
Metric . KBSAF Recall
Quantifies ABS coverage of key back-supporting atomic facts:
Metric . KBSAF Precision
Measures ABS focus on key back-supporting atomic facts:
Metric . KBSAF F1-score

### v2

# What Are They Talking About? A Benchmark of Knowledge-Grounded Discussion Summarization

@@ Abstract
Traditional dialogue summarization primarily focuses on dialogue content, assuming it comprises adequate information for a clear summary. However, this assumption often fails for discussions grounded in shared background, where participants frequently omit context and use implicit references. This results in summaries that are confusing to readers unfamiliar with the background. To address this, we introduce Knowledge-Grounded Discussion Summarization (KGDS), a novel task that produces a supplementary background summary for context and a clear opinion summary with clarified references. To facilitate research, we construct the first KGDS benchmark, featuring news-discussion pairs and expert-created multi-granularity gold annotations for evaluating sub-summaries. We also propose a novel hierarchical evaluation framework with fine-grained and interpretable metrics. Our extensive evaluation of 12 advanced large language models (LLMs) reveals that KGDS remains a significant challenge. The models frequently miss key facts and retain irrelevant ones in background summarization, and often fail to resolve implicit references in opinion summary integration.111Our benchmark is available at zhouweixiao/KGDS
@@
A Benchmark of Knowledge-Grounded Discussion Summarization
Weixiao Zhou Junnan Zhu Gengyao Li Xianfu Cheng Xinnian Liang Feifei Zhai Zhoujun Li††thanks: Corresponding Author State Key Laboratory of Complex & Critical Software Environment, Beihang University State Key Laboratory of Multimodal Artificial Intelligence Systems, Institute of Automation, CAS University of Chinese Academy of Sciences ByteDance Seed Fanyu AI Laboratory wxzhou@buaa.edu.cn junnan.zhu@nlpr.ia.ac.cn
@@ 1 Introduction
Dialogue summarization aims to distill key topics and interactions from a dialogue into a concise summary Kirstein et al. (2024); Rennard et al. (2023); Jia et al. (2023). Conventional paradigm relies primarily on dialogue content, whether in benchmark construction Gliwa et al. (2019); Chen et al. (2021); Zhu et al. (2021), methodologies Zhou et al. (2023); Tian et al. (2024); Lu et al. (2025); Zhu et al. (2025), or evaluation Wang et al. (2022); Gao and Wan (2022); Zhu et al. (2023); Tang et al. (2023, 2024b); Liu et al. (2024b); Ramprasad et al. (2024). These efforts implicitly assume that "the dialogue itself contains sufficient information to generate a clearly understandable summary for readers."
@@
However, we find this assumption has fundamental limitations and often fails, particularly when participants discuss shared background knowledge they are already familiar with. Such discussions exhibit two main traits: (1) Information Omission and Implicit Reference: Participants naturally skip mutually known details and frequently use pronouns or phrases to refer to entities or facts within the contextual background. (2) Personal Opinion: Unlike simple information interactions in general dialogues, these discussions focus on exchanging viewpoints, with participants expressing personal opinions from various perspectives. These characteristics make understanding the discussion heavily reliant on background knowledge. Consequently, traditional dialogue summarization paradigm inherits this dependency, leading to confusion for outside readers unfamiliar with the context, leaving them wondering: "What are they talking about?". We present an illustrative example in Figure 1.
@@
To address this problem, we introduce KGDS, a novel task designed to combine shared background knowledge with discussion content to create reader-centered summaries. We argue that a successful KGDS summary must achieve two complementary objectives: (1) to bridge readers’ knowledge gaps by providing the necessary background information that supports the discussion; and (2) to present readers with clear participant opinions by clarifying the implicit references within them. Thus, we formalize these requirements by modeling the task output as a Background Summary and an Opinion Summary. The background summary, which retrieves or condenses relevant background information, can be either Extractive or Abstractive, while the opinion summary integrates clarified opinions and is inherently Abstractive. In Figure 1, we provide a visual example of our task paradigms.
To advance research in this field, we construct the first KGDS benchmark, situated in a common and realistic scenario of news discussions. It consists of event-rich news articles, each paired with a human-authored discussion. To enable robust assessment, we develop multi-granularity gold evaluation components, which are annotated by experts under strict consistency controls. For the extractive background summary, we annotate coarse-grained supporting and nonsupporting paragraphs. For the abstractive background summary, we create two sets of finest-grained atomic facts Min et al. (2023), including key supporting and nonsupporting facts. For the opinion summary, we introduce clear atomic opinions as basic evaluation units that are minimized and have implicit references clarified.
Furthermore, we propose a novel hierarchical framework for evaluation. For each summarization paradigm, we first evaluate its sub-summaries separately and then aggregate their scores to derive a paradigm-level score. At the sub-summary level, our framework assesses multiple dimensions. For the two types of background summaries, we evaluate their coverage, focus, and overall quality by comparing them against annotated paragraphs or atomic facts, using paragraph index matching or LLM-based fact verification Wei et al. (2024). For the opinion summary, we assess overall quality by checking its coverage of annotated atomic opinions and identify fine-grained integration errors by classifying failure types for uncovered opinions. We also conduct human evaluation and show a strong correlation with our automatic metrics.
We evaluate 12 advanced LLMs on our benchmark under the structured-prompt Li et al. (2023) and self-reflection Shinn et al. (2023) settings. Our comprehensive results demonstrate that KGDS remains a substantial challenge, with even the top-performing models achieving an average score below 69%. We also identify several key weaknesses: LLMs struggle with the coverage-focus trade-off in background summary retrieval, frequently miss key facts during background summary generation, and fail to resolve implicit references in opinion summary integration. Furthermore, the limited effectiveness of self-reflection indicates that current LLMs lack sufficient self-correction abilities for this task. These findings highlight key bottlenecks and provide concrete directions for future advancements in coarse-grained retrieval, fine-grained generation, and knowledge integration.
@@ 2 Task Formulation
Let denote the shared background knowledge among participants and represent the discussion grounded in . The objective of KGDS is to provide a supplementary background summary and a clear opinion summary by integrating and . We define two summarization paradigms based on the type of background summary.
EBS-AOS Paradigm.
@@
Here, is the summarizer. is the EBS, defined as extractive background supporting chunks for from . Chunks are text sequences of a predefined granularity (e.g., sentences, paragraphs, etc.). is the AOS, defined as clear personal opinions of the participants with clarified implicit references.
ABS-AOS Paradigm.
@@
The definitions of and follow Eq. (1). , the ABS, is defined as abstractive background supporting information for from .
3 Benchmark Construction
@@ 3.1 Preliminary
We establish our benchmark scenario as a two-participant discussion of news content for two reasons. First, news discussions are highly prevalent in daily life, making them more representative than private scenarios such as internal meetings and medical consultations. Second, news summarization Goyal et al. (2022); Zhang et al. (2024); Liu et al. (2024a) is a well-established subfield of automatic summarization research.
@@
We collect 100 multi-domain (i.e., business, sports, and world) event-rich news articles from Google News222https://news.google.com as shared background knowledge, with a time cutoff of Oct. 2024. We preserve the original news paragraph structure and define paragraph-as-chunk as the minimum extraction granularity under the EBS-AOS paradigm.
Expert Annotators.
To ensure high-quality data, we recruit four PhD candidates specializing in NLP for our annotation tasks. Each pair of experts forms a collaborative group to conduct the full-process annotation, which includes discussion generation and the creation of multi-granularity evaluation components for background and opinion summaries.
3.2 Annotation
Human Discussion.
@@ 3.2 Annotation
This construction follows the sequence of read, understand, then discuss. For each news article, we require two experts to independently read and thoroughly comprehend its content. This preparatory step aims to ensure background consistency between the participants. Afterward, they engage in a discussion to exchange viewpoints. The entire process is open-ended, meaning the discussion initiator is chosen randomly, and the discussion topics can encompass any events, facts, or detailed information within the news article.
| Data | Scale | Avg. Tokens |
| KGDS Task Inputs | ||
| News-Discussion Pair | 100 | 729.5 |
| EBS Evaluation Components | ||
| Supporting Paragraph | 432 | 43.8 |
| Nonsupporting Paragraph | 1,005 | 42.6 |
| ABS Evaluation Components | ||
| Key Supporting Atomic Fact | 1,638 | 11.3 |
| Nonsupporting Atomic Fact | 4,996 | 11.4 |
| AOS Evaluation Components | ||
| Clear Atomic Opinion | 873 | 19.4 |
| Clarified Implicit References | 1,113 | 4.1 |
Paragraph-Level Annotation for EBS. An ideal EBS should include all paragraphs that support the discussion and exclude any that do not. To create gold labels for this, two experts independently classify each paragraph from the source news as either supporting or nonsupporting. We perform a consistency control, retaining only paragraphs where both experts agree on the label. Paragraphs with conflicting annotations are considered ambiguous and are removed from the news source to ensure data clarity. Our statistics show that out of a total of 1,696 paragraphs, 1,437 (84.7%) are annotated consistently. Table 1 provides more statistics.
Atomic-Fact-Level Annotation for ABS. Unlike coarse-grained paragraph annotation, a model-generated ABS is condensed and requires fine-grained, key-point-focused labeling. Inspired by the minimal granularity of atomic facts Liu et al. (2023b); Tang et al. (2024a), we argue that an ideal ABS should cover all key supporting facts while filtering out nonsupporting ones. Motivated by this, we create these two distinct atomic fact sets.
First, we utilize GPT-4o333Version: gpt-4o-2024-08-06 to decompose supporting paragraphs into candidate atomic facts based on principles of indivisibility, independence, and declarativity (see Appendix F.2). Two experts then independently classify each fact as either key or non-key. Through a consistency check, only facts unanimously labeled as key by both experts are considered as the gold standard, resulting in 1,638 key supporting atomic facts.
| Person1 thinks that it was wise of Rees-Zammit (him) |
| to sign with the Kansas City Chiefs (this team). |
| Person2 thinks that worrying about Jersey’s (their) |
| economic diversification is an overconcern. |
@@
Second, we use the same decomposition method to nonsupporting paragraphs to generate an initial set of nonsupporting facts. From this set, we then mask444The masking process is consistent with automatic fact verification Tang et al. (2024a), and the masked conflicting facts are not considered in ABS evaluation. any conflicting facts, defined as those that could also be inferred from supporting paragraphs. Such conflicts typically arise when identical or similar facts appear in both supporting and nonsupporting paragraphs at the atomic granularity level. For instance, two events have the same timestamp, but one is the background event while the other is not. By masking these facts, we ensure that the key supporting and nonsupporting fact sets remain non-overlapping, thereby avoiding external verification issues during evaluation. This process yields 4,996 nonsupporting atomic facts, with conflicting facts proving to be sparse at just 176 instances (3.39%).
Atomic-Opinion-Level Annotation for AOS. According to the task setup, an effective AOS must clearly present participant opinions by clarifying implicit references. To enable fine-grained evaluation, we introduce the clear atomic opinion as the basic unit, which is both minimal and has implicit references clarified. Table 2 provides two examples. We argue that the quality of an AOS can be assessed by measuring its coverage of these atomic opinions. For annotation, we first require experts to extract the main opinions from their respective utterances in the discussion. They then identify referential pronouns and phrases within these opinions and clarify them through anaphora resolution or information supplementation to produce clear opinions. Finally, the experts decompose these clear opinions into atomic opinion units, following the principles of indivisibility and independence. This annotation process is fully expert-authored. We create 873 clear atomic opinions, among which 800 contain clarified implicit references, while 73 require no clarification. Table 1 provides statistics.
4 Evaluation Framework
Our framework is comprehensive and hierarchical. For each of the two summarization paradigms (§2), we independently evaluate the sub-summaries and then aggregate their quality to assess overall performance at the paradigm level. The evaluation methods are interpretable, and the metrics are fine-grained. Figure 2 shows an example. Below, we first describe the evaluation dimensions555We do not focus on fluency and coherence because current LLMs perform well on these dimensions Song et al. (2025). at the sub-summary and paradigm levels (§4.1), and then introduce our automatic methods and metrics (§4.2).
4.1 Dimensions
Background Summary.
For the two variants of background summaries, we measure the following dimensions of system responses:
Coverage: Whether the model-extracted or -generated summary fully covers the supporting information (i.e., paragraphs or key facts).
Focus: Whether the summary focuses on supporting information while excluding nonsupporting information.